if a girl is half the age of mom and her mom is five years younger than her dad. The sum of their ages is one hundred. How old is the girl
Answer by ankor@dixie-net.com(22740)

g = girl's present age
m = Mom's age\
d = Dad's age
:
Write an equation for each statement
:
if a girl is half the age of mom
g = .5m
and her mom is five years younger than her dad.
d = m + 5
The sum of their ages is one hundred.
g + m + d = 100
Find Mom's age first, substitute .5m for g; and (m+5) for d
.5m + m + (m+5) = 100
2.5m = 100 - 5
m = 95/2.5
m = 38 yrs old is Mom
:
How old is the girl
g = .5m
g = .5(38)
g = 19 yrs
;
:
:
Check this find the sum of their ages (dad is 43, 5 yrs old than Mom)
19 + 38 + 43 = 100